# Verifying Well-Posedness of Linear PDEs using Convex Optimization

### 01 Apr 2026 v1

@@ V A PIE-Based Generation Theorem for Arbitrary $C_{0}$-Semigroups
It is well-known that any $C_{0}$-semigroup satisfies a bound of this form for some $\omega\in\mathbb{R}$ and $M\geq 1$ [2]. However, while Thm. 10 allows arbitrary growth rates $\omega$, it restricts the resulting semigroup to have gain $M=1$. In the following subsection, therefore, we establish conditions for generation of a $C_{0}$-semigroup with arbitrary gain $M$ on $L_{2}^{n}$, by testing generation of a quasicontraction semigroup in the weighted inner product space, $H=(L_{2}^{n},\left\langle,\right\rangle_{\mathcal{P}})$. We then show how these conditions can be verified via convex optimization by parameterizing $\mathcal{P}$ as a partial integral operator.
@@ V-A Generation of $C_{0}$-Semigroups on $L_{2}$
Thm. 10 provides necessary and sufficient conditions for $\{\mathcal{T},\mathcal{A}\}$ to generate a quasicontraction semigroup on a Hilbert space $H$. Applying this result directly to $H=L_{2}^{n}$, then, this only certifies generation of a $C_{0}$-semigroup with gain $M=1$. To capture more general semigroups, we instead consider the Hilbert space $H=(L_{2}^{n},\left\langle\cdot,\cdot\right\rangle_{\mathcal{P}})$ with weighted inner product $\left\langle\mathbf{u},\mathbf{v}\right\rangle_{\mathcal{P}}:=\left\langle\mathbf{u},\mathcal{P}\mathbf{v}\right\rangle_{L_{2}}$, for $\mathcal{P}:L_{2}^{n}\to L_{2}^{n}$. Here, choosing $\mathcal{P}$ to be coercive and bounded, the weighted inner product is equivalent to the standard $L_{2}$ inner product, allowing generation of a $C_{0}$-semigroup on $L_{2}^{n}$ to be verified by testing generation of a quasicontraction semigroup on $(L_{2}^{n},\left\langle\cdot,\cdot\right\rangle_{\mathcal{P}})$, yielding the following result.
@@
For given $n\in\mathbb{N}$ and $a,b]\subseteq\mathbb{R}$, let $\mathcal{T}:L_{2}^{n}[a,b]\to D$ and $\mathcal{A}:\in\mathcal{L}(L_{2}^{n}[a,b])$ be bounded linear operators, with $\mathcal{T}$ bijective onto $D$. Define $A:=\mathcal{A}\circ\mathcal{T}^{-1}:D\to L_{2}^{n}[a,b]$. For any $\omega\in\mathbb{R}$, if the following conditions are satisfied,
@@
    there exist $\epsilon_{1}>0$ and $\mathcal{P}\in\mathcal{L}(L_{2}^{n}[a,b])$ such that $\mathcal{P}\succeq\epsilon_{1}^{2}I$ and
    $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{\mathcal{P}}\leq\omega\left\langle\mathcal{T}\mathbf{v},\mathcal{T}\mathbf{v}\right\rangle_{\mathcal{P}},\qquad\forall\mathbf{v}\in L_{2}^{n}[a,b];$
@@
    $\left\lVert{(\lambda\mathcal{T}^{*}-\mathcal{A}^{*})\mathbf{v}}\right\rVert_{L_{2}}\geq\epsilon_{2}\left\lVert{\mathbf{v}}\right\rVert_{L_{2}},\qquad\forall\mathbf{v}\in L_{2}^{n}[a,b];$
then $A$ generates a $C_{0}$-semigroup on $L_{2}$ with growth rate $\omega$ and gain $M:=\frac{\sqrt{\left\lVert{\mathcal{P}}\right\rVert_{\mathcal{L}(L_{2})}}}{\epsilon_{1}}$.
@@
$\epsilon_{1}^{2}\left\lVert{\mathbf{v}}\right\rVert_{L_{2}}^{2}\leq\left\lVert{\mathbf{v}}\right\rVert_{\mathcal{P}}^{2}\leq\left\lVert{\mathcal{P}}\right\rVert_{\mathcal{L}(L_{2})}\left\lVert{\mathbf{v}}\right\rVert_{L_{2}}^{2},\quad\forall\mathbf{v}\in L_{2}^{n}.$
Defining $\epsilon:=\frac{\epsilon_{2}}{\sqrt{\left\lVert{\mathcal{P}}\right\rVert_{\mathcal{L}(L_{2})}}}$, it follows that for all $\mathbf{v}\in L_{2}^{n}$,
$\left\lVert{(\lambda\mathcal{T}^{*}-\mathcal{A}^{*})\mathbf{v}}\right\rVert_{L_{2}}^{2}\geq\epsilon_{2}^{2}\left\lVert{\mathbf{v}}\right\rVert_{L_{2}}^{2}\geq\epsilon^{2}\left\lVert{\mathbf{v}}\right\rVert_{\mathcal{P}}^{2}.$
By Thm. 10, $A$ generates a semigroup on $H:=(L_{2},\left\langle\cdot,\cdot\right\rangle_{\mathcal{P}})$ with growth rate $\omega$. This contraction semigroup is also a $C_{0}$-semigroup on $(L_{2},\left\langle\cdot,\cdot\right\rangle_{L_{2}})$, where for all $\mathbf{x}\in D$,
$\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}\leq\frac{1}{\epsilon_{1}}\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{\mathcal{P}}\leq\frac{1}{\epsilon_{1}}e^{\omega t}\left\lVert{\mathbf{x}}\right\rVert_{\mathcal{P}}\leq Me^{\omega t}\left\lVert{\mathbf{x}}\right\rVert_{L_{2}}.$
Theorem 12 provides sufficient conditions for $\{\mathcal{T},\mathcal{A}\}$ to generate a $C_{0}$-semigroup with arbitrary gain $M$ and rate $\omega$ on $L_{2}^{n}$. The condition $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{\mathcal{P}}\leq\omega\left\langle\mathcal{T}\mathbf{v},\mathcal{T}\mathbf{v}\right\rangle_{\mathcal{P}}$ herein is closely related to Lyapunov conditions for exponential stability in the PIE framework, differing only in that well-posedness does not require $\omega\leq 0$. Optimization-based methods for testing this condition have already been proposed for stability analysis in the PIE representation [14]. In the following subsection, we show how the well-posedness conditions may be similarly tested using convex optimization methods.
@@ V-B An LPI for Well-Posedness of PDEs
Having established a test for well-posedness of PIEs in Thm. 12, we now show how this result can be applied to PDEs of the form in (4), on PIE-compatible domains $D$ as in Defn. 2. For such systems, we can define an associated PIE by $\mathcal{T}$ as in Cor. 3 and $\mathcal{A}:=A\circ\mathcal{T}$ as in Eqn. (4). It has been shown that both $\mathcal{T}$ and $\mathcal{A}$ belong to the class of Partial Integral (PI) operators [14], taking the form
$(\mathcal{P}\mathbf{v})(s)$ $:=\boldsymbol{P}_{0}(s)\mathbf{v}(s)+\int_{0}^{s}\boldsymbol{P}_{1}(s,\theta)\mathbf{v}(\theta)$
$\qquad+\int_{s}^{1}\boldsymbol{P}_{2}(s,\theta)\mathbf{v}(\theta)d\theta,$
and parameterized by polynomials $\boldsymbol{P}_{0},\boldsymbol{P}_{1},\boldsymbol{P}_{2}$. Such PI operators form a class of bounded, linear operators, that it closed under adjoint and composition, with explicit formulae for these operations given in [14]. In addition, PI operator variables can be parameterized using a finite monomial basis, and positivity and equality constraints can be enforced via semidefinite programming. Using these properties, the following corollary shows how well-posedness in the PIE representation can be verified by solving a convex optimization problem on PI operator variables.
@@
For given $n,d\in\mathbb{N}$ and $B,C\in\mathbb{R}^{2n\times 4n}$, let the associated domain $D\subseteq W_{2}^{d,n}[a,b]$ as in Defn. 2 be PIE admissible. For any $\boldsymbol{A}_{i}\in\mathbb{R}^{n\times n}[s]$, define $A:D\to L_{2}^{n}[a,b]$ as in (4). Define $\mathcal{T}\in\mathcal{L}(L_{2})$ as in Cor. 3, and let $\mathcal{A}:=A\circ\mathcal{T}\in\mathcal{L}(L_{2})$. For any $\epsilon_{1},\epsilon_{2}>0$, if there exists a PI operator $\mathcal{P}$ and scalar $\omega\in\mathbb{R}$ solving the optimization program
$\mathcal{P}\succeq\epsilon_{1}^{2},$ (6)
$\mathcal{T}^{*}\mathcal{P}\mathcal{A}+\mathcal{A}^{*}\mathcal{P}\mathcal{T}\preceq 2\omega\mathcal{T}^{*}\mathcal{P}\mathcal{T},$
$([\omega+1]\mathcal{T}-\mathcal{A})([\omega+1]\mathcal{T}^{*}-\mathcal{A}^{*})\succeq\epsilon_{2}^{2}I,$
then $A$ is the infinitesimal generator of a $C_{0}$-semigroup on $L_{2}^{n}[a,b]$ with growth rate $\omega$ and gain $M:=\frac{\sqrt{\left\lVert{\mathcal{P}}\right\rVert_{\mathcal{L}(L_{2})}}}{\epsilon_{1}}$.
@@
Suppose $(\mathcal{P},\omega)$ solves the optimization program. Then, for all $\mathbf{v}\in L_{2}^{n}[a,b]$,
$\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{\mathcal{P}}=\frac{1}{2}\left\langle\mathbf{v},[\mathcal{T}^{*}\mathcal{P}\mathcal{A}+\mathcal{A}^{*}\mathcal{P}\mathcal{T}]\mathbf{v}\right\rangle_{L_{2}}\leq\omega\left\langle\mathcal{T}\mathbf{v},\mathcal{T}\mathbf{v}\right\rangle_{\mathcal{P}}.$
In addition, defining $\lambda:=\omega+1>\omega$, we obtain
$\left\lVert{(\lambda\mathcal{T}^{*}-\mathcal{A}^{*})\mathbf{v}}\right\rVert_{L_{2}}^{2}\geq\epsilon_{2}^{2}\left\lVert{\mathbf{v}}\right\rVert_{L_{2}}^{2}.$
@@
By Thm. 12, it follows that $A$ is the infinitesimal generator of a $C_{0}$-semigroup on $L_{2}^{n}[a,b]$ with growth rate $\omega$ and gain $M:=\frac{\sqrt{\left\lVert{\mathcal{P}}\right\rVert_{\mathcal{L}(L_{2})}}}{\epsilon_{1}}$. ∎
@@
By Cor. 13, well-posedness of a PIE-compatible PDE can be tested by solving the optimization program in (6). For each fixed $\omega$, this is a feasibility problem that is linear in the decision variable $\mathcal{P}$ and defined by operator inequalities on PI operators. We refer to such a program as a Linear PI Inequality (LPI). Using the PIETOOLS software suite [16], this LPI can be converted to a semidefinite program and solved with an SDP solver such as Mosek [20]. Using bijection, a smallest value of $\omega$ may then be established for which the program is feasible, thereby certifying well-posedness together with a least upper bound on the exponential growth rate. In the next section, we apply this optimization program to several PDE examples and obtain tight upper bounds on the exponential growth rate of solutions.
Well-Posedness in a different Hilbert-Space
Although Cor. 13 allows well-posedness to be verified of a broad class of PDEs, it only tests for generation of a $C_{0}$-semigroup on the Hilbert space $L_{2}^{n}[a,b]$, with standard inner product. In particular, the constraint $\mathcal{T}^{*}\mathcal{P}\mathcal{A}+\mathcal{A}^{*}\mathcal{P}\mathcal{T}\preceq 2\omega\mathcal{T}^{*}\mathcal{P}\mathcal{T}$ for $\mathcal{P}\succeq\epsilon^{2}I$ ensures that solutions satisfy $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}\leq Me^{\omega t}\left\lVert{\mathbf{x}}\right\rVert_{L_{2}}$ for some $M\geq 1$, as is necessary for $e^{tA}$ to be a $C_{0}$-semigroup on $L_{2}^{n}$, However, many interesting PDEs are not well-posed on $L_{2}^{n}$ with the standard inner product (see, e.g., Subsec. VI-C 2)). In such cases, solutions may instead satisfy a weaker bound of the form $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}\leq e^{\omega t}\left\lVert{\partial_{s}^{d}\mathbf{x}}\right\rVert_{L_{2}}$, which can be tested using non-coercive operators $\mathcal{P}\succeq\epsilon\mathcal{T}$ [21]. An open question is whether a similar modification of the conditions of Thm. 12 may be used to test well-posedness with such an alternative bound on solutions, verifying generation of a $C_{0}$-semigroup on a different Hilbert space.
VI Examples

### 02 Apr 2026 v2

@@ V A PIE-Based Well-Posedness Test for PDEs
It is well-known that any $C_{0}$-semigroup satisfies a bound of this form for some $\omega\in\mathbb{R}$ and $M\geq 1$ [2]. However, while Thm. 10 allows arbitrary growth rates $\omega$, it restricts the resulting semigroup to have gain $M=1$. In the following subsection, therefore, we establish conditions for generation of a $C_{0}$-semigroup with arbitrary gain $M$ on $L_{2}^{n}$, by testing generation of a quasicontraction semigroup in the weighted inner product space, $H=L_{2,\mathcal{P}}^{n}$. We then show how these conditions can be verified via convex optimization by parameterizing $\mathcal{P}$ as a partial integral operator.
@@ V-A Generation of $C_{0}$-Semigroups on $L_{2}$
Thm. 10 with $H=L_{2}^{n}[a,b]$ provides necessary and sufficient conditions for $\{\mathcal{T},\mathcal{A}\}$ to generate a quasicontraction semigroup on $L_{2}^{n}$. In order to verify generation of more general $C_{0}$-semigroups on $L_{2}^{n}$, we instead consider the weighted space $H=L_{2,\mathcal{P}}^{n}$, with $\mathcal{P}$ coercive and bounded. Since this weighted inner product is equivalent to the standard $L_{2}$ inner product, generation of a $C_{0}$-semigroup on $L_{2}^{n}$ can then be verified by testing generation of a quasicontraction semigroup on $L_{2,\mathcal{P}}^{n}$, leading to the following result.
@@
For $n\in\mathbb{N}$ and $a,b]\subseteq\mathbb{R}$, let $\mathcal{T}:L_{2}^{n}[a,b]\to D$ and $\mathcal{A}:\in\mathcal{L}(L_{2}^{n}[a,b])$ be bounded linear operators, with $\mathcal{T}$ bijective. Define $A:=\mathcal{A}\circ\mathcal{T}^{-1}:D\to L_{2}^{n}[a,b]$. For any $\omega\in\mathbb{R}$, if the following conditions are satisfied,
@@
    there exist $\epsilon_{1}>0$, $\mathcal{P}\in\mathcal{L}(L_{2}^{n})$ such that $\mathcal{P}\succeq\epsilon_{1}^{2}I$ and
    $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{L_{2,\mathcal{P}}}\leq\omega\left\langle\mathcal{T}\mathbf{v},\mathcal{T}\mathbf{v}\right\rangle_{L_{2,\mathcal{P}}},\quad\forall\mathbf{v}\in L_{2}^{n};$
@@
    $\left\lVert{(\lambda\mathcal{T}^{*}-\mathcal{A}^{*})\mathbf{v}}\right\rVert_{L_{2}}\geq\epsilon_{2}\left\lVert{\mathbf{v}}\right\rVert_{L_{2}},\qquad\forall\mathbf{v}\in L_{2}^{n};$
then $A$ generates a $C_{0}$-semigroup on $L_{2}^{n}$ with growth rate $\omega$ and gain $M:=\frac{\sqrt{\left\lVert{\mathcal{P}}\right\rVert_{\mathcal{L}(L_{2})}}}{\epsilon_{1}}$.
@@
$\epsilon_{1}^{2}\left\lVert{\mathbf{v}}\right\rVert_{L_{2}}^{2}\leq\left\lVert{\mathbf{v}}\right\rVert_{L_{2,\mathcal{P}}}^{2}\leq\left\lVert{\mathcal{P}}\right\rVert_{\mathcal{L}(L_{2})}\left\lVert{\mathbf{v}}\right\rVert_{L_{2}}^{2},\quad\forall\mathbf{v}\in L_{2}^{n}.$
Defining $\epsilon:=\frac{\epsilon_{2}\epsilon_{1}}{\sqrt{\left\lVert{\mathcal{P}}\right\rVert_{\mathcal{L}(L_{2})}}}$, it follows that for all $\mathbf{v}\in L_{2}^{n}$,
$\left\lVert{(\lambda\mathcal{T}^{*}-\mathcal{A}^{*})\mathbf{v}}\right\rVert_{L_{2,\mathcal{P}}}^{2}$ $\geq\epsilon_{1}^{2}\left\lVert{(\lambda\mathcal{T}^{*}-\mathcal{A}^{*})\mathbf{v}}\right\rVert_{L_{2}}^{2}$
$\qquad\geq\epsilon_{1}^{2}\epsilon_{2}^{2}\left\lVert{\mathbf{v}}\right\rVert_{L_{2}}^{2}\geq\epsilon^{2}\left\lVert{\mathbf{v}}\right\rVert_{L_{2,\mathcal{P}}}^{2}.$
By Thm. 10, $A$ generates a quasicontraction semigroup on $H:=L_{2,\mathcal{P}}^{n}$ with growth rate $\omega$. This semigroup is also a $C_{0}$-semigroup on $L_{2}^{n}$, where for all $\mathbf{x}\in D$,
$\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}\!\leq\!\frac{1}{\epsilon_{1}}\!\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2,\mathcal{P}}}\!\leq\!\frac{1}{\epsilon_{1}}e^{\omega t}\left\lVert{\mathbf{x}}\right\rVert_{L_{2,\mathcal{P}}}\!\leq\!Me^{\omega t}\left\lVert{\mathbf{x}}\right\rVert_{L_{2}}.$
Theorem 12 provides sufficient conditions for $\{\mathcal{T},\mathcal{A}\}$ to generate a $C_{0}$-semigroup with gain $M$ and rate $\omega$ on $L_{2}^{n}$. The condition $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{\mathcal{P}}\leq\omega\left\langle\mathcal{T}\mathbf{v},\mathcal{T}\mathbf{v}\right\rangle_{\mathcal{P}}$ herein is closely related to Lyapunov conditions for exponential stability in the PIE framework, differing only in that well-posedness does not require $\omega\leq 0$. Optimization-based methods for testing this condition have already been proposed for stability analysis in the PIE representation [14]. In the following subsection, we show how the well-posedness conditions may be similarly tested using convex optimization methods.
@@ V-B An LPI for Well-Posedness of PDEs
Having established a test for well-posedness of PIEs in Thm. 12, we now show how this result can be applied to PDEs of the form in Eqn. (3), on PIE-compatible domains $D$ as in Defn. 2. For such systems, we can define an associated PIE by $\mathcal{T}$ as in Cor. 3 and $\mathcal{A}:=A\circ\mathcal{T}$ as in Eqn. (6). It has been shown that both $\mathcal{T}$ and $\mathcal{A}$ belong to the class of Partial Integral (PI) operators [14], taking the form
$(\mathcal{P}\mathbf{v})(s)$ $:=\boldsymbol{P}_{0}(s)\mathbf{v}(s)+\int_{a}^{s}\boldsymbol{P}_{1}(s,\theta)\mathbf{v}(\theta)\,d\theta$
$\hskip 28.45274pt+\int_{s}^{b}\boldsymbol{P}_{2}(s,\theta)\mathbf{v}(\theta)\,d\theta,$
and parameterized by polynomials $\boldsymbol{P}_{0},\boldsymbol{P}_{1},\boldsymbol{P}_{2}$. Such PI operators form a class of bounded, linear operators, that is closed under adjoint and composition, with explicit formulae for these operations given in [14]. In addition, PI operator variables can be parameterized using a finite monomial basis, and positivity and equality constraints can be enforced via semidefinite programming. Using these properties, the following corollary shows how well-posedness in the PIE representation can be verified by solving a convex optimization problem on PI operator variables.
@@
For $n,d\in\mathbb{N}$, $a,b]\subseteq\mathbb{R}$, and $B,C\in\mathbb{R}^{nd\times nd}$, let the associated domain $D\subseteq W_{2}^{d,n}[a,b]$ as in Defn. 2 be PIE admissible. For any $\boldsymbol{A}_{i}\in\mathbb{R}^{n\times n}[s]$, define $A:D\to L_{2}^{n}[a,b]$ as in (4). Define $\mathcal{T}:L_{2}^{n}\to D$ as in Cor. 3, and let $\mathcal{A}:=A\circ\mathcal{T}\in\mathcal{L}(L_{2}^{n})$. For any $\epsilon_{1},\epsilon_{2}>0$ and $\omega\in\mathbb{R}$, if there exists a PI operator $\mathcal{P}$ and scalar $\lambda>\omega$ satisfying
$\mathcal{P}\succeq\epsilon_{1}^{2}I_{n},\qquad\mathcal{T}^{*}\mathcal{P}\mathcal{A}+\mathcal{A}^{*}\mathcal{P}\mathcal{T}$ $\preceq 2\omega\mathcal{T}^{*}\mathcal{P}\mathcal{T},$ (7)
$(\lambda\mathcal{T}-\mathcal{A})(\lambda\mathcal{T}^{*}-\mathcal{A}^{*})$ $\succeq\epsilon_{2}^{2}I_{n},$
then AA is the infinitesimal generator of a $C_{0}$-semigroup on $L_{2}^{n}[a,b]$ with growth rate $\omega$ and gain $M:=\frac{\sqrt{\left\lVert{\mathcal{P}}\right\rVert_{\mathcal{L}(L_{2})}}}{\epsilon_{1}}$.
@@
Suppose $\mathcal{P},\omega$ satisfy (7). Then, for all $\mathbf{v}\in L_{2}^{n}$,
$\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{L_{2,\mathcal{P}}}$ $=\frac{1}{2}\left\langle\mathbf{v},[\mathcal{T}^{*}\mathcal{P}\mathcal{A}+\mathcal{A}^{*}\mathcal{P}\mathcal{T}]\mathbf{v}\right\rangle_{L_{2}}$
$\leq\omega\left\langle\mathcal{T}\mathbf{v},\mathcal{T}\mathbf{v}\right\rangle_{L_{2,\mathcal{P}}},$
@@
as well as $\left\lVert{(\lambda\mathcal{T}^{*}-\mathcal{A}^{*})\mathbf{v}}\right\rVert_{L_{2}}^{2}\geq\epsilon_{2}^{2}\left\lVert{\mathbf{v}}\right\rVert_{L_{2}}^{2}$. By Thm. 12, it follows that $A$ is the infinitesimal generator of a $C_{0}$-semigroup on $L_{2}^{n}[a,b]$ with growth rate $\omega$ and gain $M$. ∎
@@
By Cor. 13, well-posedness of a PIE-compatible PDE can be tested by solving the optimization program in (7). For each fixed $\omega$, this is a feasibility problem that is linear in the decision variable $\mathcal{P}$ and defined by PI operator inequalities. We refer to such a program as a Linear PI Inequality (LPI). Using the PIETOOLS software suite [16], this LPI can be converted to a semidefinite program and solved with an SDP solver such as Mosek [21]. Using bisection, a smallest value of $\omega$ may then be established for which the program is feasible, thereby certifying well-posedness together with a least upper bound on the exponential growth rate. In the next section, we apply this optimization program to several PDE examples and obtain tight upper bounds on the exponential growth rate of solutions.
VI Illustrations of the Well-Posedness Test